The total surface area of a right circular cone is one and one-half times the lateral area. The radius of the base is equal to _______inches if the slant height is 6 inches.
Answer by mananth(16946)

Total surface area of cone = pir^2+pirl = pir(r+l)
Lateral surface area = pi r l
Given condition
pir(r+l) = 3/2 * pi r l
(r+l) = 3/2* l

2r+2l = 3l
l=2r
6 = 2r
r=3 inches